OpenStudy (anonymous):

what two number added together =-6 ; but when multiplied = 210?

OpenStudy (anonymous):

i mean when added together = -37

OpenStudy (anonymous):

so; what two #'s added together = -36 but when multiplied together = 210?

OpenStudy (shamil98):

x + y = -36 xy = 210 x = -y - 36 -y(y-36) = 210 simplify and solve as a quadratic.

OpenStudy (anonymous):

-30 and -7

OpenStudy (shamil98):

-37? or -36 ..

OpenStudy (anonymous):

my bad again.. -37